OpenStudy (anonymous):

A satellite is orbiting the earth at a height where the gravitational pull of the earth is given by the constant, 9.2 . The force (in N) exerted by the earth on the satellite is 9.2 times the mass of the satellite (in kilograms). Write a function to define the force as a function of the mass. If m is mass and F is force, what is the function?

OpenStudy (jamesj):

You there?

OpenStudy (anonymous):

Yea. I'm here!

OpenStudy (jamesj):

So in other words, you're told that the Force is 9.2 times the mass. For example, if the satellite weighed 1 kg, the force would be F = 9.2(1kg) = 9.2 N What if m = 1000 kg, well then F = 9.2(1000) = 920 N The question is asking you to generalize this now. Suppose the mass is m. What is the force?

OpenStudy (anonymous):

BRB

OpenStudy (anonymous):

F(m) = 9.2 F(m) = 9.2 + m F(m) = 9.2m F(m)= 9.2/m ?

OpenStudy (jamesj):

Look at my two example above. Which equation in your options fits them?

OpenStudy (anonymous):

F(m)=9.2m?

OpenStudy (jamesj):

Yes, obviously. And it makes complete sense. By Newton's Second law, F = ma that is, the force F is proportional to mass m

OpenStudy (anonymous):

Oh yes. Thank you JamesJ